147 NSR.2 You are the technical consultant for a car-chase se- quence in an action movie. In a certain part of the scene, the director wants a car to round a certain curve while braking from 66 mi/h to rest before it travels 290 ft along the curve. You measure the r dius of the curve to be 590 ft. Is this scene possible? Defend your response and suggest an alternative scene if it is not possible. a- dvanced Equations N8.24 through N8.26 in problem N85.13 alots speed in its circular 8A.1 THE ANSWER IS NO, YOU CAN CHANGE THE INITIAL SPEED TO 22 M/S OR YOU CAN CHANGE THE GROUND TO CONCRETE WHICH HAS A COEFFICIENT OF FRICTION OF .6-.85, AS OPPOSED TO ASPHALT, WHICH IS .9
Answer #1

r Axis of rotation f frequency in revolutions/s

using the formula W^2=W^2_o+2\alpha \theta , where \alpha=angular acceleration and \theta= angle until get rest

solving for \alpha and putting the numerical data, knowing that 1m=3ft, it follows that

\alpha =-\frac{W^2_o}{2\theta }=-\frac{\frac{V}{R}^2}{2\frac{S}{R} }=-\frac{\frac{22m/s^2}{196,6m}}{2\frac{96.6m}{196.6m}}=-2.50m/s^2,where S= distance travels until get rest, R=radius V=velocity

this is a typical acceleration for a car, therefore, is possible